# Debunking the CUDA Myth Towards GPU-based AI Systems : Evaluation of the Performance and Programmability of Intel’s Gaudi NPU for AI Model Serving

### v1

@@ Abstract.
With the rise of AI, NVIDIA GPUs have become the de facto standard for AI system design. This paper presents a comprehensive evaluation of Intel Gaudi NPUs as an alternative to NVIDIA GPUs for AI model serving. First, we create a suite of microbenchmarks to compare Intel Gaudi-2 with NVIDIA A100, showing that Gaudi-2 achieves competitive performance not only in primitive AI compute, memory, and communication operations but also in executing several important AI workloads end-to-end. We then assess Gaudi NPU’s programmability by discussing several software-level optimization strategies to employ for implementing critical FBGEMM operators and vLLM, evaluating their efficiency against GPU-optimized counterparts. Results indicate that Gaudi-2 achieves energy efficiency comparable to A100, though there are notable areas for improvement in terms of software maturity. Overall, we conclude that, with effective integration into high-level AI frameworks, Gaudi NPUs could challenge NVIDIA GPU’s dominance in the AI server market, though further improvements are necessary to fully compete with NVIDIA’s robust software ecosystem.
@@ 1. Introduction
(Performance) To enable detailed experimental studies and analyses, we first develop microbenchmarks targeting Gaudi NPUs to stress-test their ability to maximize performance in several key compute, memory, and communication primitives. In this work, we use Intel’s second-generation Gaudi NPU (Gaudi-2) and NVIDIA’s A100 GPU as comparison points111The hardware and software architecture of Intel’s recently announced Gaudi-3 is virtually identical to that of Gaudi-2 (but with limited availability), except that Gaudi-3 offers higher compute and memory throughput, thanks to its chiplet-based design., as both processors are manufactured using TSMC’s 7nm technology node and are supported by an HBM2E-based memory subsystem, providing comparable performance (Table 1). Our microbenchmarking revealed that Gaudi-2 demonstrates highly competitive performance in important primitive AI operations, particularly for tasks involving regular compute and memory accesses. However, Gaudi-2 did fall short of A100 in certain scenarios involving fine-grained data accesses and collective communications across a small number of processors.
After confirming Gaudi NPU’s competitiveness against NVIDIA GPUs in primitive AI operations through our microbenchmark-based analysis, we next evaluate both systems at the end-to-end AI application level. Specifically, we focus on recommendation systems (RecSys) and large language models (LLMs), as these two are among the most widely deployed AI models in cloud environments. Our analysis of end-to-end AI applications revealed that Gaudi-2 achieves competitive energy efficiency for RecSys and superior energy efficiency for LLMs compared to the A100.
@@ 2.2. Intel Gaudi Software Architecture
(Programming model) The TPC programming model is based on the Single Program Multiple Data (SPMD) parallel programming paradigm. In this model, the target workload is partitioned across different TPCs, all of which execute the same TPC program. The workload distribution is performed by partitioning the index space (equivalent to CUDA grid), enabling each TPC to process different data independently. The index space can be divided up to five dimensions, and each member of the index space is allocated with an indivisible unit of work processed by a single TPC.
@@
Figure 2 illustrates how a single TPC operates to conduct a simple element-wise vector addition operation (pseudo-code shown in Figure 3). A TPC program is typically structured using a for-loop, which iteratively executes vector-wide “LoadComputeStore” operations. For optimal performance, the TPC programmers are advised to carefully partition the index space across the TPCs (Figure 3(a), line 25-27) while maximizing the performance of individual TPCs (Figure 3(b)). Two important best practices are recommended for TPC programmers. First, to maximize memory bandwidth utilization, the TPC’s data access granularity should be aligned to 256 bytes as this is the minimum access granularity for global memory (e.g., a single step in the depth dimension in Figure 2 and line 18 in Figure 3(b), which determines the memory access granularity, is sized at 256-bytes). Second, to fully utilize the TPC processor, it is recommended that programmers manually unroll the for-loop to maximize both instruction-level and memory-level parallelism (e.g., a single step in the width dimension, which determines the magnitude of loop unrolling and thus the number of parallel operations, is sized at 4. See line 40-63 in Figure 3(b)). This recommendation stems from the fact that TPC instructions have an average architectural latency of 4 processor cycles (i.e., the effect of executing a TPC instruction is reflected in the architectural state 4 cycles later) (Intel, 2024b). By unrolling four TPC instructions within a given iteration of a for-loop, the processor pipeline can be better utilized and memory access latency can be better hidden.
@@
(Graph compiler) Intel’s Gaudi NPU comes with a software suite called Intel Gaudi SDK (Intel, 2024e), which is tightly integrated with PyTorch and TensorFlow. This integration allows developers to utilize familiar tools while taking advantage of Gaudi NPU’s hardware acceleration. Gaudi SDK includes a graph compiler that converts AI models into a format optimized for execution on Gaudi NPUs. The graph compiler not only applies high-level model optimizations, such as operator and kernel fusion, but it also performs hardware-specific optimizations. These include MME configuration adjustments for high MME utilization (Figure 1, further discussed in Section 3.2) and operator pipelining between the MME and TPC. When an MME operation is followed by a TPC operation (e.g., GEMM followed by an activation function), the graph compiler breaks them into smaller, independent sub-operations to enable pipelined execution. This approach helps hide latency and reduce overall execution time. The graph compiler fully controls the process of orchestrating data transfers between the MME and TPC for pipelined execution, using on-chip shared memory.
@@
3. Characterizing Gaudi NPU Performance
3.1. Motivation and Evaluation Methodology
@@
(Methodology) All experiments discussed in the rest of this paper are conducted using an HLS-Gaudi-2 server (which contains eight Gaudi-2 chips connected via RoCE) and a DGX A100 server (which contains eight A100 GPUs connected via NVSwitch and NVLink) (Table 1). On the software side, we use Intel Gaudi Software v1.18.0, which is based on PyTorch 2.4, along with the TPC-C SDK for custom Gaudi kernel development. For the GPU system evaluation, we use PyTorch 2.4 and CUDA 12.4 for GPU kernel development. All experimental results presented in this paper assume the BF16 data type, except when we evaluate end-to-end RecSys models which utilize FP32222State-of-the-art RecSys model serving is done using FBGEMM (Khudia et al., 2021) which is optimized for FP32 without supporting BF16, so our evaluation utilizes FP32 for RecSys serving. . The non-RecSys evaluation results for FP32 and their key takeaways were practically identical to those for BF16, so we omit presenting those results for brevity.
For our microbenchmark analysis targeting primitive AI operations (Section 3.2 to Section 3.4), we focus on absolute performance and its resource utilization. For end-to-end AI workload analysis (Section 3.5), we evaluate both performance and energy-efficiency. When evaluating the energy-efficiency of end-to-end AI workloads, each system’s power consumption is measured using nvidia-smi (NVIDIA, 2024b) for A100 and hl-smi (Intel, 2024) for Gaudi-2.
@@ 3.2. Primivite “Compute” Operations
We first show how a single TPC’s performance is improved by applying the two aforementioned best practices. In Figure 8(a), we measure the compute throughput by varying our microbenchmarks’ data access granularity from 2 to 2,048 bytes, without loop unrolling. The results clearly show the significant performance drop when the data access granularity is set lower than 256 bytes, which is Gaudi’s minimum memory access granularity. At data access granularities higher than 256 bytes, the overall throughput saturates at around 55 GFLOPS for TRIAD and around 30 GFLOPS for SCALE and ADD. From this point, we explore how much further performance improvements can be achieved by unrolling the for-loop in Algorithm 1 (discussed in Section 2.2). As shown in Figure 8(b), the compute throughput of SCALE improves remarkably, while ADD and TRIAD achieve only slight improvements as the loop unrolling factor increases. Overall, these results highlight the importance of loop unrolling in TPC-C kernels to exploit instruction-level and memory-level parallelism for high-performance.
@@
Consequently, to explore how much Gaudi-2 and A100 can saturate the peak compute throughput of their vector engines (11 TFLOPS with Gaudi TPCs and 39 TFLOPS with A100 SIMD Cores), the experiments in Figure 8(d,e,f) artificially increase the operational intensity of ADD, SCALE, and TRIAD. Specifically, starting from the default STREAM benchmark configuration, we gradually increase the operational intensity (defined as the number of operations performed relative to the number of memory accesses) up to the point where the achieved compute throughput becomes saturated. At lower operational intensities, the three microbenchmarks become memory-bound, so Gaudi-2 demonstrates slightly higher compute throughput than A100 thanks to its higher memory bandwidth (Table 1). At higher operational intensities where the workloads become compute-bound, A100 achieves much higher compute throughput because of its higher vector computation power. In terms of compute utilization, Gaudi-2’s compute throughput saturates at approximately 5.5 TFLOPS, 5.5 TFLOPS, and 10.9 TFLOPS for the ADD, SCALE, and TRIAD operations, reaching , , and of its maximum 11 TFLOPS throughput, respectively. In comparison, A100’s compute throughput saturates at around 19.4 TFLOPS, 19.4 TFLOPS, and 38.2 TFLOPS for the ADD, SCALE, and TRIAD operations, similarly reaching , , and of its maximum 39 TFLOPS throughput, respectively.
@@ 3.5. End-to-End Application-level Analysis
(Single-device model serving) At the time of this writing, the embedding layer implemented in Intel’s Gaudi SDK (Intel, 2024e) achieved, on average, 37% of the performance of its GPU counterpart in the RecSys configurations we explore in this work. We implemented a custom version of the embedding layer using TPC-C, which incorporates several key performance optimization strategies employed in the GPU-optimized FBGEMM library (Section 4.1 details our implementation strategy), and we utilize it as Gaudi’s backend kernel library for all our evaluations in end-to-end RecSys model serving.
@@
In Figure 11, we show Gaudi-2’s end-to-end speedup (top) and energy-efficiency (bottom) over A100 for the two RecSys models, RM1 and RM2. Overall, Gaudi-2 experiences an average performance degradation of 22% and 18% for RM1 and RM2, respectively. While Gaudi-2 achieves higher performance with wide embedding vectors and large batch sizes (maximum 1.36 speedup), which can be attributed to Gaudi-2’s higher compute throughput (Table 1), it generally falls short of A100 for all other model configurations. For instance, in the memory-intensive RM2, Gaudi-2 exhibits a significant performance drop in embedding vector sizes less than 256-byte, regardless of batch size (a maximum of 70% performance loss). This is primarily due to Gaudi-2’s 256-byte minimum memory access granularity and reduced memory bandwidth utilization for small number of vector gather operations, as highlighted in Section 3.3. In terms of power consumption, Gaudi-2 consumed an average of 11% less absolute power than A100 across all evaluated data points in RM1 and RM2. This result was counterintuitive to our expectations, given that Gaudi-2’s TDP is 50% higher than the A100 (Table 1), which may be coming from the differences in the microarchitecture designs of the two processors and how they implement DVFS mechanisms. Overall, due to an increase in average latency with Gaudi-2, its end-to-end energy consumption is increased by 8% than A100 on average across all our evaluation points in RM1 and RM2.
@@
As for LLMs, the leftmost heatmap in Figure 12(a) shows Gaudi-2’s speedup over A100 when serving Llama-3.1-8B-Instruct over a single device. Across all batch sizes and output token lengths, Gaudi-2 consistently outperforms A100, achieving an average speedup of 1.47 (with a maximum 1.70 speedup). This performance advantage generally stems from Gaudi-2’s high peak FLOPS and memory bandwidth, benefiting both the compute-bound prefill stage and the memory-bound decoding stage. Although Gaudi-2’s theoretical performance advantage over A100 is 1.4 in GEMM throughput and 1.2 in memory bandwidth (Table 1), Gaudi-2 achieves an even greater speedup due to its superior compute utilization across various GEMM shapes, as discussed in Section 3.2. The leftmost heatmap in Figure 12(b) shows a single Gaudi-2’s energy-efficiency improvement over A100. In general, Gaudi-2 exhibited lower power consumption than A100 for small batch sizes, while its power consumption increased with larger batches and eventually surpassed A100 at the highest batch sizes. On average, Gaudi-2 exhibited 91% absolute power consumption of A100 across all data points. As discussed in Figure 7(a), for small GEMM shapes, Gaudi-2 activates only a subset of its large MME units so a possible explanation for Gaudi-2’s lower power consumption over small batches is that it more aggressively power-gates its circuitry via DVFS. Overall, Gaudi-2’s higher absolute performance and comparable power consumption yield an average 64% higher energy-efficiency than A100.
@@
(Multi-device model serving) The three rightmost heatmaps in Figure 12(a) illustrate Gaudi-2’s speedup when serving the Llama-3.1-70B-Instruct model across 2, 4, and 8 devices using tensor-parallelism (TP) (Shoeybi et al., 2019). The multi-device LLM serving results generally align with the trends observed in the single-device Llama-3.1-8B-Instruct serving, with average speedups of 1.29, 1.32, and 1.35 for 2, 4, and 8 devices, respectively. Interestingly, the speedup level increases with the number of devices used, attributed to the collective communication performance provided by Intel’s HLS-Gaudi-2 server. As discussed in Section 3.4, the use of P2P links for communication makes the performance of all-reduce (the collective communication primitive in tensor-parallelism) to be proportionally higher to the number of devices involved in collective communication, enabling higher system-level performance as more devices are employed for multi-device LLM serving. Energy-efficiency trends, illustrated in the three rightmost heatmaps in Figure 12(b), generally align with those observed in single-device serving. Across all configurations using 2, 4, or 8 devices, Gaudi-2 consistently demonstrates lower average power consumption, consuming around 79% of the power of A100. Overall, Gaudi-2 achieves energy efficiency improvements of 1.66, 1.75, and 1.80 over A100 for 2, 4, and 8 devices.

### v2

@@ Abstract.
This paper presents a comprehensive evaluation of Intel Gaudi NPUs as an alternative to NVIDIA GPUs, which is currently the de facto standard in AI system design. First, we create a suite of microbenchmarks to compare Intel Gaudi-2 with NVIDIA A100, showing that Gaudi-2 achieves competitive performance not only in primitive AI compute, memory, and communication operations but also in executing several important AI workloads end-to-end. We then assess Gaudi NPU’s programmability by discussing several software-level optimization strategies to employ for implementing critical FBGEMM operators and vLLM, evaluating their efficiency against GPU-optimized counterparts. Results indicate that Gaudi-2 achieves energy efficiency comparable to A100, though there are notable areas for improvement in terms of software maturity. Overall, we conclude that, with effective integration into high-level AI frameworks, Gaudi NPUs could challenge NVIDIA GPU’s dominance in the AI server market, though further improvements are necessary to fully compete with NVIDIA’s robust software ecosystem.
@@ 1. Introduction
(Performance) To enable detailed experimental studies and analyses, we first develop a set of microbenchmarks targeting Gaudi NPUs to stress-test their ability to maximize performance in several key compute, memory, and communication primitives. In this work, we use Intel’s second-generation Gaudi NPU (Gaudi-2) and NVIDIA’s A100 GPU as comparison points111The hardware and software architecture of Intel’s recently announced Gaudi-3 is virtually identical to that of Gaudi-2 (but with limited availability), except that Gaudi-3 offers higher compute and memory throughput, thanks to its chiplet-based design., as both processors are manufactured using TSMC’s 7nm technology node and are supported by an HBM2E-based memory subsystem, providing comparable performance (Table 1). Our microbenchmarking revealed that Gaudi-2 demonstrates highly competitive performance in important primitive AI operations, particularly for tasks involving regular compute and memory accesses. However, Gaudi-2 did fall short of A100 in certain scenarios involving fine-grained data accesses and collective communications across a small number of processors.
After confirming Gaudi NPU’s competitiveness against NVIDIA GPUs in primitive AI operations through our microbenchmark-based analysis, we next evaluate both systems at the end-to-end AI application level. Specifically, we focus on recommendation systems (RecSys) and large language models (LLMs), as these two are among the most widely deployed AI models in cloud environments. Our analysis of end-to-end AI applications revealed that Gaudi-2 achieves 28% lower energy-efficiency for RecSys but 50% higher energy-efficiency for LLMs compared to the A100.
@@ 2.2. Intel Gaudi Software Architecture
(Programming model) The GPU programming system follows the Single Instruction Multiple Thread (SIMT) model, which falls under the Single Program Multiple Data (SPMD) paradigm. To support efficient SIMT execution, GPUs include unique microarchitectural support, such as a large register file for fine-grained, massive multithreading (Gebhart et al., 2011), dynamic branch divergence resolution (Fung et al., 2007; Rhu and Erez, 2013; Shoushtary et al., 2024), and warp-wide memory coalescing (Chatterjee et al., 2014; Meng et al., 2010). In contrast, Gaudi lacks these features and instead uses a single-threaded programming model optimized for data-level rather than thread-level parallelism.
In NVIDIA’s CUDA, programmers can leverage the WMMA (Warp Matrix Multiply and Accumulate) APIs (Markidis et al., 2018) to directly utilize Tensor Cores alongside conventional SIMD Cores for computation within low-level CUDA kernels. Furthermore, these operations can be extended to a higher-level PyTorch API (line 10 of Figure 2(a) and line 9 of Figure 2(b)). In contrast, the Gaudi SDK currently restricts direct access to the MME units, allowing programmers to explicitly control only the TPCs. Instead, access to MME units is limited to the PyTorch level (Figure 2(a), line 14), thereby constraining programmers from directly optimizing performance involving both MME and TPC at a lower level. This limitation can be addressed through the Gaudi graph compiler’s optimization pass, which we will discuss later. In Gaudi’s TPC programming model, the target workload is partitioned across different TPCs, which execute the same TPC program. Workload distribution is performed by partitioning the index space (equivalent to CUDA grid), enabling each TPC to process different data independently. The index space can be divided up to five dimensions, and each member of the index space is allocated with an indivisible unit of work processed by a single TPC.
@@
Figure 3 illustrates how a single TPC operates to conduct a simple element-wise vector addition operation (pseudo-code shown in Figure 2(c)). A TPC program is typically structured using a for-loop, which iteratively executes vector-wide “LoadComputeStore” operations. For optimal performance, the TPC programmers are advised to carefully partition the index space across the TPCs while maximizing the performance of individual TPCs. Two important best practices are recommended for TPC programmers. First, to maximize memory bandwidth utilization, the TPC’s data access granularity should be aligned to 256 bytes as this is the minimum access granularity for global memory (e.g., line 7 in Figure 2(c) and a single step in the depth dimension in Figure 3, which determines the memory access granularity, is sized at 256-bytes). Second, to fully utilize the TPC processor, it is recommended that programmers manually unroll the for-loop to maximize both instruction-level and memory-level parallelism (e.g., a single step in the width dimension, which determines the magnitude of loop unrolling and thus the number of parallel operations, is sized at 4). This recommendation stems from the fact that TPC instructions have an average architectural latency of 4 processor cycles (i.e., the effect of executing a TPC instruction is reflected in the architectural state 4 cycles later) (Intel, 2024b). By unrolling four TPC instructions within a given iteration of a for-loop, the processor pipeline can be better utilized and memory access latency can be better hidden. To reduce the burden of manual loop unrolling, the TPC programming system provides preprocessor directives such as “#pragma unroll” (line 16 in Figure 2(c)).
@@
(Graph compiler) Intel’s Gaudi NPU comes with a software suite called Intel Gaudi SDK (Intel, 2024f), which is tightly integrated with PyTorch and TensorFlow. This integration allows developers to utilize familiar tools while taking advantage of Gaudi NPU’s hardware acceleration. Gaudi SDK includes a graph compiler that converts AI models into a format optimized for execution on Gaudi NPUs. The graph compiler applies high-level model optimizations, such as operator and kernel fusion. For example, an MLIR (Lattner et al., 2021)-based operation fuser selects arbitrary subgraphs of element-wise, reduction, and normalization operations, then JIT-fuses and compiles them into TPC kernels (Bobba et al., 2024; Kaplan, 2024; Intel, 2024e). This improves performance by reducing memory bandwidth usage and enabling tensor shape-aware optimizations, unlike shape-agnostic kernel libraries. The graph compiler also performs hardware-specific optimizations, which include MME configuration adjustments for high MME utilization (Figure 1, further discussed in Section 3.2) and operator pipelining between MME and TPC. When an MME operation is followed by a TPC operation (e.g., GEMM followed by an activation function), the graph compiler breaks them into smaller, independent sub-operations to enable pipelined execution. This approach helps hide latency and reduce overall execution time. The graph compiler controls the process of orchestrating data transfers between the MME and TPC for pipelined execution, using on-chip shared memory as an intermediate buffer to minimize redundant off-chip memory accesses.
@@
3. Characterizing Gaudi NPU Performance
3.1. Motivation and Evaluation Methodology
@@
(Methodology) All experiments discussed in the rest of this paper are conducted using an HLS-Gaudi-2 server (which contains eight Gaudi-2 chips connected via RoCE) and a DGX A100 server (which contains eight A100 GPUs connected via NVSwitch and NVLink) (Table 1). On the software side, we use Intel Gaudi Software v1.18.0, which is based on PyTorch 2.4, along with the TPC-C SDK for custom Gaudi kernel development. For the GPU system evaluation, we use PyTorch 2.4 and CUDA 12.4 for GPU kernel development. All experimental results presented in this paper assume the BF16 data type, except when we evaluate end-to-end RecSys models which utilize FP32. The non-RecSys evaluation results for FP32 and their key takeaways were practically identical to those for BF16, so we omit presenting those results for brevity.
For our microbenchmark analysis targeting primitive AI operations (Section 3.2 to Section 3.4), we focus on absolute performance and its resource utilization. When evaluating GEMM operations, we used PyTorch API, which defaults to cuBLAS (NVIDIA, 2024a) as its backend library, choosing the most optimal GEMM implementation per target device. For end-to-end AI workload analysis (Section 3.5), we evaluate both performance and energy-efficiency. When evaluating the energy-efficiency of end-to-end AI workloads, each system’s power consumption is measured using nvidia-smi (NVIDIA, 2024b) for A100 and hl-smi (Intel, 2024) for Gaudi-2.
@@ 3.2. Primivite “Compute” Operations
We first show how a single TPC’s performance is improved by applying the two aforementioned best practices. In Figure 8(a), we measure the compute throughput by varying our microbenchmarks’ data access granularity from 2 to 2,048 bytes, without loop unrolling. The results clearly show the significant performance drop when the data access granularity is set lower than 256 bytes, which is Gaudi’s minimum memory access granularity. At data access granularities higher than 256 bytes, the overall throughput saturates at around 55 GFLOPS for TRIAD and around 30 GFLOPS for SCALE and ADD. From this point, we explore how much further performance improvements can be achieved by unrolling the for-loop in Algorithm 1 (discussed in Section 2.2). As shown in Figure 8(b), the compute throughput of SCALE improves remarkably, while ADD and TRIAD achieve only slight improvements as the loop unrolling factor increases. Both TRIAD and ADD load vectors from two arrays, resulting in two load instructions and one compute instruction (v_bf16_mac_b for TRIAD and v_bf16_add_b for ADD) for each loop iteration. In contrast, SCALE requires only one load instruction and one compute instruction (v_bf16_mul_b) from a single array, for each loop iteration, providing more pipeline opportunities to hide the 4 TPC processor cycle latency and benefiting more from loop unrolling. Overall, these results highlight the importance of loop unrolling in TPC-C kernels to exploit instruction-level and memory-level parallelism for high-performance.
@@
Consequently, to explore how much Gaudi-2 and A100 can saturate the peak compute throughput of their vector engines (11 TFLOPS with Gaudi TPCs and 39 TFLOPS with A100 SIMD Cores), the experiments in Figure 8(d,e,f) artificially increase the operational intensity of ADD, SCALE, and TRIAD. Specifically, starting from the default STREAM benchmark configuration, we gradually increase the operational intensity (defined as the number of operations performed relative to the number of memory accesses) up to the point where the achieved compute throughput becomes saturated. At lower operational intensities, the three microbenchmarks become memory-bound, so Gaudi-2 demonstrates slightly higher compute throughput than A100 thanks to its higher memory bandwidth (Table 1). At higher operational intensities where the workloads become compute-bound, A100 achieves much higher compute throughput because of its higher vector computation power. In terms of compute utilization, Gaudi-2’s compute throughput saturates at approximately 5.5 TFLOPS, 5.5 TFLOPS, and 10.9 TFLOPS for the ADD, SCALE, and TRIAD operations, reaching , , and of its maximum 11 TFLOPS throughput, respectively. In comparison, A100’s compute throughput saturates at around 19.4 TFLOPS, 19.4 TFLOPS, and 38.2 TFLOPS for the ADD, SCALE, and TRIAD operations, similarly reaching , , and of its maximum 39 TFLOPS throughput, respectively (Figure 8(d,e,f)).
@@ 3.5. End-to-End Application-level Analysis
(Single-device model serving) At the time of this writing, the embedding layer implemented in Intel’s Gaudi SDK (Intel, 2024f) achieved, on average, 37% of the performance of its GPU counterpart in the RecSys configurations we explore in this work. To assess the potential of Gaudi-2’s hardware architecture without being constrained by the limitations of the Gaudi SDK’s current implementations, we designed a custom version of the embedding layer using TPC-C. This design incorporates several key performance optimization strategies employed in the GPU-optimized FBGEMM library (Section 4.1 details our implementation strategy), and we utilize it as Gaudi’s backend kernel library for all our evaluations in end-to-end RecSys model serving.
@@
In Figure 11, we show Gaudi-2’s end-to-end speedup (top) and energy-efficiency (bottom) over A100 for the two RecSys models, RM1 and RM2. Overall, Gaudi-2 experiences an average performance degradation of 22% and 18% for RM1 and RM2, respectively. While Gaudi-2 achieves higher performance with wide embedding vectors and large batch sizes (maximum 1.36 speedup), which can be attributed to Gaudi-2’s higher compute throughput (Table 1), it generally falls short of A100 for all other model configurations. For instance, in the memory-intensive RM2, Gaudi-2 exhibits a significant performance drop in embedding vector sizes less than 256-byte, regardless of batch size (a maximum of 70% performance loss). This is primarily due to Gaudi-2’s 256-byte minimum memory access granularity and reduced memory bandwidth utilization for small number of vector gather operations, as highlighted in Section 3.3. In terms of power consumption, Gaudi-2 consumed an average of 12% more absolute power than A100 in RM1 and RM2. This result aligns with our expectations, given that Gaudi’s TDP is 50% higher than A100 (Table 1). Overall, with an increase in average latency with Gaudi-2, its end-to-end energy consumption is increased by an average 28% than A100 for RM1 and RM2.
@@
As for LLMs, the leftmost heatmap in Figure 12(a) shows Gaudi-2’s speedup over A100 when serving Llama-3.1-8B-Instruct over a single device. Across all batch sizes and output token lengths, Gaudi-2 consistently outperforms A100, achieving an average speedup of 1.47 (with a maximum 1.70 speedup). This performance advantage generally stems from Gaudi-2’s high peak FLOPS and memory bandwidth, benefiting both the compute-bound prefill stage and the memory-bound decoding stage (Figure 12(b)). Although Gaudi-2’s theoretical performance advantage over A100 is 1.4 in GEMM throughput and 1.2 in memory bandwidth (Table 1), Gaudi-2 achieves an even greater speedup due to its superior compute utilization across various GEMM shapes, as discussed in Section 3.2. The leftmost heatmap in Figure 13 shows a single Gaudi-2’s energy-efficiency improvement over A100. In general, Gaudi-2 exhibited lower power consumption than A100 for small batch sizes, while its power consumption increased with larger batches and eventually surpassed A100 at the highest batch sizes. On average, Gaudi-2 exhibited only an average 1% higher power consumption than A100, despite its 50% higher TDP. As discussed in Figure 7(a), for small GEMM shapes, Gaudi-2 activates only a subset of its large MME units so a possible explanation for Gaudi-2’s lower power consumption over small batches is that it more aggressively power-gates its circuitry via DVFS. Overall, Gaudi-2’s higher absolute performance and comparable power consumption yield an average 48% higher energy-efficiency than A100 for single-device LLM serving.
@@
(Multi-device model serving) The three rightmost heatmaps in Figure 12(a) illustrate Gaudi-2’s speedup when serving the Llama-3.1-70B-Instruct model across 2, 4, and 8 devices using tensor-parallelism (TP) (Shoeybi et al., 2019). The multi-device LLM serving results generally align with the trends observed in the single-device Llama-3.1-8B-Instruct serving, with average speedups of 1.29, 1.32, and 1.35 for 2, 4, and 8 devices, respectively. Interestingly, the speedup level increases with the number of devices used, attributed to the collective communication performance provided by Intel’s HLS-Gaudi-2 server. As discussed in Section 3.4, the use of P2P links for communication makes the performance of all-reduce (the collective communication primitive in tensor-parallelism) to be proportionally higher to the number of devices involved in collective communication, enabling higher system-level performance as more devices are employed for multi-device LLM serving. Energy-efficiency trends, illustrated in the three rightmost heatmaps in Figure 13, generally align with those observed in single-device serving. Across all multi-device LLM servings, Gaudi-2 consistently demonstrates lower average power consumption, consuming around 88% of the power of A100. Overall, Gaudi-2 achieves energy-efficiency improvements of 1.48, 1.51, and 1.56 over A100 for 2, 4, and 8 devices.